OpenStudy (anonymous):

using the completed-square form (x-4)^2-21 to solve the equation x^2-8x-5=0, leaving the answer in exact (surd) form

OpenStudy (anonymous):

\[x^2-8x=5\] \[(x-4)^2=5+4^2=21\] \[x-4=\pm\sqrt{21}\] \[x=4\pm\sqrt{21}\]

OpenStudy (anonymous):

i love the word "surd" we don't use it much anymore

OpenStudy (anonymous):

you can use quadratic formula also